# ADBM: Adversarial Diffusion Bridge Model for Reliable Adversarial Purification

### v1

@@ 4.1 Training Objective
For detailed derivations, please refer to Sec. B.1.
Following Eq. 3, ADBM uses ${\mathbf{x}}_{\theta}({\mathbf{x}}_{t}^{d},t):=\frac{1}{\sqrt{\bar{\alpha}_{t}}}({\mathbf{x}}_{t}^{d}-\sqrt{1-\bar{\alpha}_{t}}{\boldsymbol{\epsilon}}_{\theta}({\mathbf{x}}_{t}^{d},t))$ to approximate the clean example ${\mathbf{x}}_{0}$. Thus the loss of ADBM can be computed by
$L$ $=\mathbb{E}_{{\boldsymbol{\epsilon}},t}\|{\mathbf{x}}_{0}-{\mathbf{x}}_{\theta}({\mathbf{x}}_{t}^{d},t)\|^{2}=\mathbb{E}_{{\boldsymbol{\epsilon}},t}\left\|\frac{k_{t}-\sqrt{\bar{\alpha}_{t}}}{\sqrt{\bar{\alpha}_{t}}}{\boldsymbol{\epsilon}}_{a}-\frac{\sqrt{1-\bar{\alpha}_{t}}}{\sqrt{\bar{\alpha}_{t}}}({\boldsymbol{\epsilon}}-{\boldsymbol{\epsilon}}_{\theta}({\mathbf{x}}_{t}^{d},t))\right\|^{2}$ (8)
$=\mathbb{E}_{{\boldsymbol{\epsilon}},t}\left[\frac{1-\bar{\alpha}_{t}}{\bar{\alpha}_{t}}\left\|\frac{\bar{\alpha}_{T}\sqrt{1-\bar{\alpha}_{t}}}{(1-\bar{\alpha}_{T})\sqrt{\bar{\alpha}_{t}}}{\boldsymbol{\epsilon}}_{a}+{\boldsymbol{\epsilon}}-{\boldsymbol{\epsilon}}_{\theta}({\mathbf{x}}_{t}^{d},t)\right\|^{2}\right].$
Omitting $\frac{1-\bar{\alpha}_{t}}{\sqrt{\bar{\alpha}_{t}}}$ as in Ho et al. [15], for a given ${\mathbf{x}}_{0}$, the final loss $L_{b}$ of ADBM is given by:
$L_{b}=$ $\mathbb{E}_{{\boldsymbol{\epsilon}},t}\left[\left\|\frac{\bar{\alpha}_{T}\sqrt{1-\bar{\alpha}_{t}}}{(1-\bar{\alpha}_{T})\sqrt{\bar{\alpha}_{t}}}{\boldsymbol{\epsilon}}_{a}+{\boldsymbol{\epsilon}}-{\boldsymbol{\epsilon}}_{\theta}({\mathbf{x}}_{t}^{d},t)\right\|^{2}\right],$ (9)
where ${\mathbf{x}}_{t}^{d}=\sqrt{\bar{\alpha}_{t}}{\mathbf{x}}_{0}+\sqrt{1-\bar{\alpha}_{t}}{\boldsymbol{\epsilon}}+\frac{\bar{\alpha}_{T}(1-\bar{\alpha}_{t})}{\sqrt{\bar{\alpha}_{t}}(1-\bar{\alpha}_{T})}{\boldsymbol{\epsilon}}_{a}$. The training process of ADBM is detailed in the blue block of Figure 2.
Comparing Eq. 4 and Eq. 9, $L_{d}$ and $L_{b}$ are quite similar except that $L_{b}$ has two additional scaled ${\boldsymbol{\epsilon}}_{a}$ in both the input and prediction objective of ${\boldsymbol{\epsilon}}_{\theta}$. Thus in practice, the training of ADBM can fine-tune the pre-trained diffusion checkpoint with $L_{b}$, avoiding training from scratch. As $t$ decreases to 0, the coefficient of ${\boldsymbol{\epsilon}}_{a}$ diminishes. This complies with the intuition that ${\boldsymbol{\epsilon}}_{a}$ is gradually eliminated.
Figure 2: The illustration of ADBM training. The blue block represents the training objective, and the green block represents the extra module for adversarial noise generation. Black arrows denote the computation of $L_{b}$, blue arrows denote the computation of $L_{c}$ used for generating adversarial noise, and red arrows denote the direction of the gradient flow when calculating the adversarial noise.
@@ 4.2 Adversarial Noise Generation
To better align with this goal, we propose to generate adversarial noise with the help of the classifier. During ADBM training, we input $\hat{{\mathbf{x}}}_{0}$ into the classifier, where $\hat{{\mathbf{x}}}_{0}=\frac{1}{\sqrt{\bar{\alpha}_{t}}}({\mathbf{x}}_{t}^{d}-\sqrt{1-\bar{\alpha}_{t}}{\boldsymbol{\epsilon}}_{\theta}({\mathbf{x}}_{t}^{d},t))$. The classification loss $L_{c}$ is then given by $L_{c}(f_{\theta_{c}}(\hat{{\mathbf{x}}}_{0}),y)$, where $y$ denotes the category label of ${\mathbf{x}}_{0}$ and $f_{\theta_{c}}$ denotes the classifier. Finally, we compute $\frac{\partial L_{c}}{\partial{\boldsymbol{\epsilon}}_{a}}$ to obtain ${\boldsymbol{\epsilon}}_{a}$ that maximizes $L_{c}$. Figure 2 illustrates this process with red arrows indicating the gradient flow direction for maximizing $L_{c}$. Note that the parameters of the classifier are kept fixed during this process, ensuring that ADBM retains the plug-and-play functionality (without modifying the model to be protected) of AP methods.
Any gradient-based attack method generally can be used for maximizing $L_{c}$, but in practice, some particular proxy attack should be specified. Based on previous experiences, the PGD attack is a popular proxy that has a good generalization ability against various attacks during inference [5, 32, 7], and thus we also used such proxy attack in the ADBM training. Our experiments also validate that this proxy can generalize to other attacks such as AutoAttack, C&W [33], and DeepFool [34] attacks. Note that during the training of ADBM, both $t$ and ${\boldsymbol{\epsilon}}$ in Eq. 9 are randomly sampled for each optimization step. This introduces a significant increase in computational cost when maximizing $L_{c}$ with PGD (EOT should be used). To address this issue, we propose an alternative method: instead of random sampling, we eliminate the randomness within a specific optimization step. In other words, for a given optimization step, we sample $t$ and ${\boldsymbol{\epsilon}}$ once and then fix their values throughout the entire training step (including generating ${\boldsymbol{\epsilon}}_{a}$ using methods like PGD and optimizing ${\boldsymbol{\epsilon}}_{\theta}$). This approach avoids the time-consuming EOT computation and ensures the calculation of adversarial noise that is truly effective for ADBM training. Note that we only eliminate randomness in the ADBM training. During inference, the randomness is kept, as discussed next.
@@ 4.4 Theoretical Analysis
Given an adversarial example ${\mathbf{x}}_{0}^{a}$ and assuming the training loss $L_{b}\leq\delta$, the distance between the purified example of ADBM and the clean example ${\mathbf{x}}_{0}$, denoted as $\|\hat{{\mathbf{x}}}_{0}-{\mathbf{x}}_{0}\|$, is bounded by $\delta$ (constant omitted) in expectation when using a one-step DDIM sampler. Specifically, we have $\mathbb{E}_{{\boldsymbol{\epsilon}}}\left[\|\hat{{\mathbf{x}}}_{0}-{\mathbf{x}}_{0}\|^{2}\right]\leq\frac{(1-\bar{\alpha}_{T})T}{{\bar{\alpha}_{T}}}\delta$, where $\frac{(1-\bar{\alpha}_{T})T}{{\bar{\alpha}_{T}}}$ is the constant.
@@
Please see the full proof in Sec. B.2. ∎
Theorem 1 implies that if the training loss of ADBM converges to zero, it can perfectly remove adversarial noises by employing a one-step DDIM sampler. While for DiffPure, we cannot derive such strong theoretical guarantee (The bound provided in Theorem 3.2 of Nie et al. [16] is larger than $\|\epsilon_{a}\|$ and thus cannot be zero). Moreover, the subsequent theorem demonstrates the superiority of ADBM over DiffPure.
@@
(sketch) Eq. 10 and Eq. 11 are derived using Bayes’ rule, where $p({\mathbf{x}}_{0}|\hat{{\mathbf{x}}}_{t})\propto p({\mathbf{x}}_{0})p(\hat{{\mathbf{x}}}_{t}|{\mathbf{x}}_{0})$. From the perspective of SDE, if timestep is infinite, $\{{\mathbf{x}}_{t}\}_{t:0\to T}$ and $\{\hat{{\mathbf{x}}}_{t}\}_{t:T\to 0}$ follow the same distribution [23]. And given that $k_{t}<\sqrt{\bar{\alpha}}_{t}$ for any $1\leq t\leq T$, the inequality $P(B)>P(D)$ always holds. Please see the full proof in Sec. B.3. ∎
@@ 5.1 Experimental Settings
Datasets and network architectures. We conducted comprehensive experiments on popular datasets, including SVHN [35], CIFAR-10 [21], and Tiny-ImageNet [36], together with a large-scale dataset ImageNet-100 [37]. [^2] We cannot afford to conduct experiments on the full ImageNet-1K [38]. As an alternative, we used ImageNet-100, a curated subset of ImageNet-1K featuring 100 categories with the original resolution of ImageNet-1K. All these datasets consist of RGB images, whose resolution is $32\times 32$ for SVHN and CIFAR-10, $64\times 64$ for Tiny-ImageNet, and $224\times 224$ for ImageNet-100. We adopted the widely used WideResNet-28-10 (WRN-28-10), WRN-70-16, WRN-28-10, and ResNet-50 [39] architectures as classifiers on SVHN, CIFAR-10, Tiny-ImageNet, and ImageNet-100, respectively. As for the diffusion models, we employed the UNet architecture [40] improved by Song et al. [23], specifically, the DDPM++ continuous variant. Pre-trained diffusion checkpoints are required for DiffPure. We directly used the checkpoint provided by Song et al. [23] for CIFAR-10 and we used their code to train the checkpoints for other datasets. These trained checkpoints were used in DiffPure and served as baselines for ADBM.
Fine-tuning settings of ADBM. The adversarial noise was computed in the popular norm-ball setting $\|{\boldsymbol{\epsilon}}_{a}\|_{\infty}\leq 8/255$. When computing ${\boldsymbol{\epsilon}}_{a}$, we used PGD with three iteration steps and a step size of $8/255$. Other settings followed the standard configuration used in Song et al. [23]. The fine-tuning steps were set to 30K, which is about 1/10 the training steps of the original diffusion models. In each fine-tuning step, the value of $T$ in Eq. 9 was uniformly sampled from 100 to 200. Note that when fine-tuning the diffusion models, the parameters of the classifier were kept frozen. Additional settings are provided in Sec. C.1.
Defense configurations of ADBM. Unless otherwise specified, the forward diffusion steps were set to be 100 for SVHN and CIFAR-10 and 150 for Tiny-ImageNet and ImageNet-100, respectively. The reverse steps were set to be five. The reverse process used a DDIM sampler. These configurations were also applied to DiffPure for a fair comparison.
@@
WRN-70-16 [41] AT $91.10$ $65.92$ $8.26$ $27.56$ $33.91$
WRN-70-16 [42] $88.54$ $64.26$ $12.06$ $32.29$ $36.20$
WRN-70-16 Augment w/ Diff [24] $88.74$ $66.18$ $9.76$ $28.73$ $34.89$
WRN-70-16 Augment w/ Diff [25] $93.25$ 70.72 $8.48$ $28.98$ $36.06$
MLP+WRN-28-10 [13] AP $91.89$ $4.56$ $8.68$ $7.25$ $6.83$
UNet+WRN-70-16 [14] $87.93$ $37.65$ $36.87$ $57.81$ $44.11$
UNet+WRN-70-16 DiffPure+Guide [26] $93.16$ $22.07$ $28.71$ $35.74$ $28.84$
UNet+WRN-70-16 Diff+ScoreOpt [28] $91.41$ $13.28$ $10.94$ $28.91$ $17.71$
UNet+WRN-70-16 DiffPure+Langevin [27] $92.18$ $43.75$ $39.84$ $55.47$ $46.35$
UNet+WRN-70-16 DiffPure [16] $92.5\pm 0.5$ $42.2\pm 2.1$ $44.3\pm 1.3$ $60.8\pm 2.3$ $49.1\pm 1.7$

### v3

@@ 4.1 Training Objective
For detailed derivations, please refer to Sec. C.1.
Following Eq. 3, ADBM uses ${\mathbf{x}}_{\theta}({\mathbf{x}}_{t}^{d},t):=\frac{1}{\sqrt{\bar{\alpha}_{t}}}({\mathbf{x}}_{t}^{d}-\sqrt{1-\bar{\alpha}_{t}}{\bm{\epsilon}}_{\theta}({\mathbf{x}}_{t}^{d},t))$ to approximate the clean example ${\mathbf{x}}_{0}$. Thus the loss of ADBM can be computed by
$L$ $=\mathbb{E}_{{\bm{\epsilon}},t}\|{\mathbf{x}}_{0}-{\mathbf{x}}_{\theta}({\mathbf{x}}_{t}^{d},t)\|^{2}=\mathbb{E}_{{\bm{\epsilon}},t}\left\|\frac{k_{t}-\sqrt{\bar{\alpha}_{t}}}{\sqrt{\bar{\alpha}_{t}}}{\bm{\epsilon}}_{a}-\frac{\sqrt{1-\bar{\alpha}_{t}}}{\sqrt{\bar{\alpha}_{t}}}({\bm{\epsilon}}-{\bm{\epsilon}}_{\theta}({\mathbf{x}}_{t}^{d},t))\right\|^{2}$ (8)
$=\mathbb{E}_{{\bm{\epsilon}},t}\left[\frac{1-\bar{\alpha}_{t}}{\bar{\alpha}_{t}}\left\|\frac{\bar{\alpha}_{T}\sqrt{1-\bar{\alpha}_{t}}}{(1-\bar{\alpha}_{T})\sqrt{\bar{\alpha}_{t}}}{\bm{\epsilon}}_{a}+{\bm{\epsilon}}-{\bm{\epsilon}}_{\theta}({\mathbf{x}}_{t}^{d},t)\right\|^{2}\right].$
Omitting $\frac{1-\bar{\alpha}_{t}}{\sqrt{\bar{\alpha}_{t}}}$ as in Ho et al. (2020), for a given ${\mathbf{x}}_{0}$, the final loss $L_{b}$ of ADBM is given by:
$L_{b}=$ $\mathbb{E}_{{\bm{\epsilon}},t}\left[\left\|\frac{\bar{\alpha}_{T}\sqrt{1-\bar{\alpha}_{t}}}{(1-\bar{\alpha}_{T})\sqrt{\bar{\alpha}_{t}}}{\bm{\epsilon}}_{a}+{\bm{\epsilon}}-{\bm{\epsilon}}_{\theta}({\mathbf{x}}_{t}^{d},t)\right\|^{2}\right],$ (9)
where ${\mathbf{x}}_{t}^{d}=\sqrt{\bar{\alpha}_{t}}{\mathbf{x}}_{0}+\sqrt{1-\bar{\alpha}_{t}}{\bm{\epsilon}}+\frac{\bar{\alpha}_{T}(1-\bar{\alpha}_{t})}{\sqrt{\bar{\alpha}_{t}}(1-\bar{\alpha}_{T})}{\bm{\epsilon}}_{a}$. The training process of ADBM is detailed in the blue block of Figure 2.
Comparing Eq. 4 and Eq. 9, $L_{d}$ and $L_{b}$ are quite similar except that $L_{b}$ has two additional scaled ${\bm{\epsilon}}_{a}$ in both the input and prediction objective of ${\bm{\epsilon}}_{\theta}$. Thus in practice, the training of ADBM can fine-tune the pre-trained diffusion checkpoint with $L_{b}$, avoiding training from scratch. As $t$ decreases to 0, the coefficient of ${\bm{\epsilon}}_{a}$ diminishes. This complies with the intuition that ${\bm{\epsilon}}_{a}$ is gradually eliminated.
Figure 2: The illustration of ADBM training. The blue block represents the training objective, and the green block represents the extra module for adversarial noise generation. Black arrows denote the computation of $L_{b}$, blue arrows denote the computation of $L_{c}$ used for generating adversarial noise, and red arrows denote the direction of the gradient flow when calculating ${\bm{\epsilon}}_{a}$.
@@ 4.2 Adversarial Noise Generation
To better align with this goal, we propose to generate adversarial noise with the help of the classifier. During ADBM training, we input $\hat{{\mathbf{x}}}_{0}$ into the classifier, where $\hat{{\mathbf{x}}}_{0}=\frac{1}{\sqrt{\bar{\alpha}_{t}}}({\mathbf{x}}_{t}^{d}-\sqrt{1-\bar{\alpha}_{t}}{\bm{\epsilon}}_{\theta}({\mathbf{x}}_{t}^{d},t))$. The classification loss $L_{c}$ is then given by $L_{c}(f_{\theta_{c}}(\hat{{\mathbf{x}}}_{0}),y)$, where $y$ denotes the category label of ${\mathbf{x}}_{0}$ and $f_{\theta_{c}}$ denotes the classifier. Finally, we compute $\frac{\partial L_{c}}{\partial{\bm{\epsilon}}_{a}}$ to obtain ${\bm{\epsilon}}_{a}$ that maximizes $L_{c}$. Figure 2 illustrates this process with red arrows indicating the gradient flow direction for maximizing $L_{c}$. The process of first maximizing $L_{c}$ then minimizing $L_{b}$ shares a similar procedure with adversarial training, but we note that: 1) The parameters of the classifier should be kept fixed throughout this process, ensuring that ADBM retains the plug-and-play functionality (without modifying the model to be protected) of AP methods; 2) The introduction of $\epsilon_{a}$ should align with Eq. 9 to effectively address the trade-off in DiffPure between noise removal and data recovery. In addition, although Lin et al. (2024) designed an impressive AT method named AToP for various AP methods, they claim that AToP is ineffective for diffusion models, highlighting the challenges of enhancing diffusion-based purification models.
Any gradient-based attack method generally can be used for maximizing $L_{c}$, but in practice, some particular proxy attack should be specified. Based on previous experiences, the PGD attack is a popular proxy that has a good generalization ability against various attacks during inference (Athalye et al., 2018; Croce & Hein, 2020; Madry et al., 2018), and thus we also used such proxy attack in the ADBM training. Our experiments also validate that this proxy can generalize to other attacks such as AutoAttack, C&W (Carlini & Wagner, 2017), and DeepFool (Moosavi-Dezfooli et al., 2016) attacks. Note that during the training of ADBM, both $t$ and ${\bm{\epsilon}}$ in Eq. 9 are randomly sampled for each optimization step. This introduces a significant increase in computational cost when maximizing $L_{c}$ with PGD (EOT should be used). To address this issue, we propose an alternative method: instead of random sampling, we eliminate the randomness within a specific optimization step. In other words, for a given optimization step, we sample $t$ and ${\bm{\epsilon}}$ once and then fix their values throughout the entire training step (including generating ${\bm{\epsilon}}_{a}$ using methods like PGD and optimizing ${\bm{\epsilon}}_{\theta}$). The insight behind our approach is that maintaining randomness makes the maximization of $L_{c}$ via PGD challenging without EOT, and eliminating the randomness within a specific optimization step avoids the time-consuming EOT computation and ensures the calculation of adversarial noise that is truly effective for ADBM training. Note that we only eliminate randomness in the ADBM training. During inference, the randomness is kept, as discussed next.
@@ 4.4 Theoretical Analysis
Given an adversarial example ${\mathbf{x}}_{0}^{a}$ and assuming the training loss $L_{b}\leq\delta$, the distance between the purified example of ADBM and the clean example ${\mathbf{x}}_{0}$, denoted as $\|\hat{{\mathbf{x}}}_{0}-{\mathbf{x}}_{0}\|$, is bounded by $\delta$ (constant omitted) in expectation when using a one-step DDIM sampler. Specifically, we have $\mathbb{E}_{{\bm{\epsilon}}}\left[\|\hat{{\mathbf{x}}}_{0}-{\mathbf{x}}_{0}\|^{2}\right]\leq\frac{(1-\bar{\alpha}_{T})T}{{\bar{\alpha}_{T}}}\delta$, where $\frac{(1-\bar{\alpha}_{T})T}{{\bar{\alpha}_{T}}}$ is the constant.
@@
Please see the full proof in Sec. C.2. ∎
Theorem 1 implies that if the training loss of ADBM converges to zero, it can perfectly remove adversarial noises by employing a one-step DDIM sampler. While for DiffPure, we cannot derive such strong theoretical guarantee (The bound provided in Theorem 3.2 of Nie et al. (2022) is larger than $\|\epsilon_{a}\|$ and thus cannot be zero). Moreover, the subsequent theorem demonstrates the superiority of ADBM over DiffPure.
@@
(sketch) Eq. 10 and Eq. 11 are derived using Bayes’ rule, where $p({\mathbf{x}}_{0}|\hat{{\mathbf{x}}}_{t})\propto p({\mathbf{x}}_{0})p(\hat{{\mathbf{x}}}_{t}|{\mathbf{x}}_{0})$. From the perspective of SDE, if timestep is infinite, $\{{\mathbf{x}}_{t}\}_{t:0\to T}$ and $\{\hat{{\mathbf{x}}}_{t}\}_{t:T\to 0}$ follow the same distribution (Song et al., 2021). And given that $k_{t}<\sqrt{\bar{\alpha}}_{t}$ for any $1\leq t\leq T$, the inequality $P(B)>P(D)$ always holds. Please see the full proof in Sec. C.3. ∎
@@ 5.1 Experimental Settings
Datasets and network architectures. We conducted comprehensive experiments on popular datasets, including SVHN (Netzer et al., 2011), CIFAR-10 (Krizhevsky et al., 2009), and Tiny-ImageNet (Le & Yang, 2015), together with a large-scale dataset ImageNet-100  [^2] https://www.kaggle.com/datasets/ambityga/imagenet100. We cannot afford to conduct experiments on the full ImageNet-1K (Russakovsky et al., 2015). As an alternative, we used ImageNet-100, a curated subset of ImageNet-1K featuring 100 categories with the original resolution of ImageNet-1K. All these datasets consist of RGB images, whose resolution is $32\times 32$ for SVHN and CIFAR-10, $64\times 64$ for Tiny-ImageNet, and $224\times 224$ for ImageNet-100. We adopted the widely used WideResNet-28-10 (WRN-28-10), WRN-70-16, WRN-28-10, and ResNet-50 (He et al., 2016) architectures as classifiers on SVHN, CIFAR-10, Tiny-ImageNet, and ImageNet-100, respectively. As for the diffusion models, we employed the UNet architecture (Ronneberger et al., 2015) improved by Song et al. (2021), specifically, the DDPM++ continuous variant. Pre-trained diffusion checkpoints are required for DiffPure. We directly used the checkpoint provided by Song et al. (2021) for CIFAR-10 and we used their code to train the checkpoints for other datasets. These trained checkpoints were used in DiffPure and served as baselines for ADBM.
Fine-tuning settings of ADBM. The adversarial noise was computed in the popular norm-ball setting $\|{\bm{\epsilon}}_{a}\|_{\infty}\leq 8/255$. When computing ${\bm{\epsilon}}_{a}$, we used PGD with three iteration steps and a step size of $8/255$. Other settings followed the standard configuration used in Song et al. (2021). The fine-tuning steps were set to 30K, which is about 1/10 the training steps of the original diffusion models. In each fine-tuning step, the value of $T$ in Eq. 9 was uniformly sampled from 100 to 200. Note that when fine-tuning the diffusion models, the parameters of the classifier were kept frozen. Additional settings are provided in Sec. D.1.
Defense configurations of ADBM. Unless otherwise specified, the forward diffusion steps were set to be 100 for SVHN and CIFAR-10 and 150 for Tiny-ImageNet and ImageNet-100, respectively. The reverse sampling steps were set to be five. The reverse process used a DDIM sampler. These configurations were also applied to DiffPure for a fair comparison.
@@
WRN-70-16 (Gowal et al., 2020) AT $91.10$ $65.92$ $8.26$ $27.56$ $33.91$
WRN-70-16 (Rebuffi et al., 2021) $88.54$ $64.26$ $12.06$ $32.29$ $36.20$
WRN-70-16 Augment w/ Diff (Gowal et al., 2021) $88.74$ $66.18$ $9.76$ $28.73$ $34.89$
WRN-70-16 Augment w/ Diff (Wang et al., 2023) $93.25$ 70.72 $8.48$ $28.98$ $36.06$
MLP+WRN-28-10 (Shi et al., 2021) AP $91.89$ $4.56$ $8.68$ $7.25$ $6.83$
UNet+WRN-70-16 (Yoon et al., 2021) $87.93$ $37.65$ $36.87$ $57.81$ $44.11$
UNet+WRN-70-16 DiffPure+Guide (Wang et al., 2022) $93.16$ $22.07$ $28.71$ $35.74$ $28.84$
UNet+WRN-70-16 Diff+ScoreOpt (Zhang et al., 2024) $91.41$ $13.28$ $10.94$ $28.91$ $17.71$
UNet+WRN-70-16 DiffPure+Langevin (Zhang et al., 2023) $92.18$ $43.75$ $39.84$ $55.47$ $46.35$
UNet+WRN-70-16 DiffPure (Nie et al., 2022) $92.5\pm 0.5$ $42.2\pm 2.1$ $44.3\pm 1.3$ $60.8\pm 2.3$ $49.1\pm 1.7$